OpenStudy (anonymous):

PLEASE HELP URGENT ! what is the 7th of the geometric sequence where a1 = 3 and a5 = -768

OpenStudy (amistre64):

nothings THAT urgent :/

OpenStudy (amistre64):

we should start by remembering the general formula for a geometric sequence

OpenStudy (amistre64):

an = a1*r^(n-1)

OpenStudy (amistre64):

we know a1, and we know the value when n=5; from this we can grab the r value

OpenStudy (amistre64):

can you fill those in for me?

OpenStudy (anonymous):

kind of

OpenStudy (amistre64):

give it a try, ill know if its right or not :)

OpenStudy (anonymous):

the answer is going to be -3,072

OpenStudy (amistre64):

it might be, but lets see if we can build up to that

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

lets fill in what we are given: a1 = 3 and a5 = -768 when n=5 an = a1 * r^(n-1) a5 = -786 = 3 * r^(5-1) -786/3 = r^(4) -292 = r^(4) , 4rt the whole thing to find r 4rt(-292) = r as is this is not a possible sequence to me. an even root can be negative across the reals

OpenStudy (amistre64):

an even root canNOT .... i passed typing class, honest :)

OpenStudy (amistre64):

check your problem and make sure you havent got an error in the post

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

i got a feeling the a1 = -3 for this to work like you want it to

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

it is

OpenStudy (amistre64):

4rt(292) = r then this is what we get for r take the general formula again, knowing this r and n=7 a7 = -3 * (4rt(292))^(6) and a calculator should popout your intended results

OpenStudy (anonymous):

can it be 12,288

OpenStudy (amistre64):

no

OpenStudy (amistre64):

...can what be?

OpenStudy (amistre64):

a7 is going to be a negative value

OpenStudy (anonymous):

-12,288

OpenStudy (amistre64):

adjust for a mismath and i get: a7 = -3 * 256^(6/4) = -12288 , so yes

OpenStudy (anonymous):

can you help me with thi one ? What is the sum of the geometric sequence -4, 28, -196 ?

OpenStudy (amistre64):

why did you think that the answer "is going to be -3,072"?

OpenStudy (anonymous):

i hd my ngatives wrng

OpenStudy (anonymous):

signs

OpenStudy (amistre64):

ah, itll prolly be best to post up a new question with that sum in it; i have to go for now